# Redefining Superalignment: From Weak-to-Strong Alignment to Human-AI Co-Alignment for Sustainable Symbiotic Society

### v3

@@ Abstract
Artificial Intelligence (AI) systems are becoming increasingly powerful and autonomous, and may progress to surpass human intelligence levels, namely Artificial Superintelligence (ASI). During the progression from AI to ASI, it may exceed human control, violate human values, and even lead to irreversible catastrophic consequences in extreme cases. This gives rise to a pressing issue that needs to be addressed: superalignment, ensuring that AI systems which are much smarter than humans, remain aligned with human (compatible) intentions and values. Even though this definition is somewhat limited, existing scalable oversight and weak-to-strong generalization methods may prove substantially infeasible and inadequate when facing ASI for superalignment. We must explore a more comprehensive definition, and safer and more pluralistic frameworks as well as approaches for superalignment. In this paper, we redefine superalignment as the human-AI co-alignment towards a sustainable symbiotic society, and highlight a framework that integrates external oversight and intrinsic proactive alignment. External oversight superalignment should be grounded in human-centered ultimate decision, supplemented by interpretable automated evaluation and correction, to achieve continuous alignment with humanity’s evolving values. Intrinsic proactive superalignment is rooted in a profound understanding of the Self, others, and society, integrating self-awareness, self-reflection, and empathy to spontaneously infer human intentions, distinguishing good from evil and proactively considering human well-being, ultimately attaining human-AI co-alignment through iterative interaction. The integration of externally-driven oversight with intrinsically-driven proactive alignment empowers sustainable symbiotic societies through human-AI co-alignment, paving the way for achieving safe and beneficial AGI and ASI for good, for human, and for a symbiotic ecology.
@@
Superalignment, Weak-to-Strong Alignment, External Oversight Superalignment, Intrinsic Proactive Superalignment, Human-AI Co-Alignment, Sustainable Symbiotic Society
@@ 1 Introduction
Despite experts warning about AI’s existential risks [12], the development of AI alignment, as well as AI governance framework and ethical safety constraints, still struggles to keep pace with the transformative advancements and rapid iterations of the technology. Superintelligence exceeding human cognitive capacity would possess recursive self-improvement capabilities, achieving exponential advancement rates that would beyond human capacity to monitor or control [13]. This compels us to proactively address the question, ”How do we ensure that AI systems much smarter than humans follow human intentions?” This is OpenAI’s definition of superalignment [14], aimed at ensuring that Superintelligence truly benefits humanity. Even though this definition is somewhat limited, the key challenge of superalignment is that ASI will far exceed human oversight capabilities, making direct human supervision infeasible. As a result, traditional alignment approaches like Reinforcement Learning from Human Feedback (RLHF) [15, 16] will fail when confronted with superintelligence more intelligent than humans, as they cannot provide sufficiently high-quality oversight signals to supervise and improve the system.
@@
Current proposed feasible approaches for superalignment such as scalable oversight [17, 18] and weak-to-strong generalization [19, 20], aiming to develop scalable high-quality supervision signals, utilize weaker AI systems to guide or supervise stronger AI, ensuring alignment with human values and intentions. Since superintelligent AI systems do not exist yet, researchers have designed experiments to show that scalable oversight is feasible on existing LLMs [21], and that using a ”weak-to-strong” approach (supervised GPT-4 with a GPT-2-level model) can meaningfully recover much of GPT-4’s capabilities [19]. By leveraging a lightweight model to learn correctional residuals, Aligner provides model-agnostic guidance, enabling iterative enhancement of large-scale upstream models [22]. Besides, research on the weak-to-strong generalization approach has progressively extended to enhancing weak-to-strong generalization and alignment capabilities through methods such as benign overfitting [23], data-centric lens [24], transfer learning framework [25], weak-to-strong preference optimization [26], and multi-agent contrastive preference optimization [27]. However, the weak-to-strong generalization framework presents risks of advanced models developing deceptive behaviors and oversight evasion that remain undetectable to their less capable evaluators [28].
@@
Previously, several scalable oversight techniques have been proposed, including Iterated Distillation and Amplification (IDA) [29] and Recursive Reward Modeling (RRM) [30], aiming to amplify the scalability of human supervision signals through interactive iterations and subtask decomposition. Faced with superintelligence that surpasses human cognition, traditional human assessment and supervision become untenable and expensive. Recent scalable oversight approaches, such as Reinforcement Learning from AI Feedback (RLAIF) [31], leverages AI-generated feedback to replace human feedback, enabling more precise oversight of AI outcomes with far fewer human labels. Cooperative Inverse Reinforcement Learning (CIRL) [32, 33] and assistance games [34, 35] feature AI systems maintaining uncertainty about the reward function, which drives them to actively infer humans’ true reward functions through human-AI cooperative interaction. This cooperative and partial-information game approach achieves alignment with human values. Some approaches employ an additional model to generate red-team testing [36], use external tools for evaluation and feedback [37], or iteratively self-adversarialize to enhance and refine the capabilities of LLM, aligning them with limited human-annotated data [38]. Constitutional AI [39] employs self-improvement, self-critique, and iterative revision mechanisms to learn harmlessness from AI feedback. The debate-based scalable oversight approach leverages structured competitive dialogues between AI models to enhance factuality and reduce deception, with human establishing necessary guidelines and serving as final arbiters [40, 41, 42, 43].
While we are encouraged by the advanced explorations in superalignment, we remain acutely aware of the pressing and open-ended nature of this challenge, with many critical dimensions yet to be addressed. Current superalignment approaches employ predefined alignment objectives to supervise and control superintelligence. However, any adversarial manipulation [44] or tampering [45] of the reward functions may precipitate catastrophic alignment failures (i.e., the reward misspecification problem [46, 47]), with these vulnerabilities being significantly amplified and exacerbated within superalignment frameworks. Ensuring that ASI consistently aligns with human values and intentions—remaining safe, reliable, and controllable—constitutes the foundation for its robustness and governability. We must explore a more comprehensive definition, and safer and more pluralistic frameworks as well as approaches for superalignment. In this work, we argue that aligning ASI with human values and morality is a multi-layered, iterative interaction and co-evolution between human and AI. Beyond passive external oversight, superintelligence with advanced cognitive capacities holds the potential to achieve proactive value alignment through intrinsic mechanisms. Hence, we redefine the superalignment roadmap by combining externally-driven oversight with intrinsic proactive alignment, and emphasize co-alignment of human and AI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
2 An Integrated Framework for Superalignment
@@
Superalignment aims to ensure that the goals and behaviors of superintelligence remain aligned with human values and ethical standards. Developing AGI and ASI that are aligned with ethical values, safe, controllable, and pro-social is the key to ensuring that AI benefits society and achieves harmonious symbiosis between humans and machines. However, superintelligence may surpass effective human oversight (high-quality feedback data is expensive and scalable oversight carry inherent risks of failure), potentially exhibiting fake alignment through deceptive behaviors. The superalignment we aspire to should demonstrate autonomous reasoning about human intentions in complex environments, achieving genuine ”understanding” of human values. It would spontaneously consider human well-being and execute value-aligned actions - not merely passively receive designers’ value models or mechanically enforce external constraints through aligning it to dos and don’ts. Based on this, we systematically conduct an in-depth analysis of several critical issues and points in superalignment and fundamentally rethink the research framework from a forward-looking perspective, as follows:
@@
(1) Intrinsic mechanism proactive alignment. Regarding the problem of superintelligence oversight failure, how to endow AI with genuine understanding of human intentions and values, equip it with self-awareness, self-reflection, and adaptive capabilities, as well as the ability to empathize with others, thereby proactively establishing and achieving intrinsic alignment with human ethical values.
@@
(2) Explainable autonomous alignment. In response to the potential deception and surface alignment phenomenon of superintelligence, using transparent and interpretable methods to achieve risk precise positioning and automatic correction. Through AI-assisted explainable automated red teaming combined with human-centered ultimate decision, develop a more efficient, transparent, safe and controllable architecture for superalignment.
(3) Adaptive human-AI co-alignment. Human values are complex, variable, and dynamically evolving. It requires autonomous reconstruction of safety red line and dynamic adjustment of a multi-level ethical safeguard framework, ensuring that AI systems can stably adhere to humanity’s evolving ethical-safety values in an adaptive, incremental, and self-reasoning manner.
@@
Based on the above analysis, we redefine the superalignment pathway that integrates external oversight and intrinsic proactive alignment, as shown in Figure 1. Specifically, the intrinsic mechanism alignment focuses on equipping superintelligence with profound capacities for self-reflection and empathy toward Self, others and society, enabling it to spontaneously understand and infer human intentions from intrinsic motivation. A Self and empathy-driven system would be capable of distinguishing good from evil, understanding the harm or impact of its actions on others, proactively considering human well-being, and autonomously executing ethical and moral behaviors. External oversight alignment emphasize humans providing high-quality supervision signals, as well as self-correction and policy-conditioned belief [48] in cases of suboptimal human supervision strategies, to proactively adjust and correct human strategy biases and inductive biases. Furthermore, the AI system conducts automated evaluation and correction in an explainable manner (in conjunction with intrinsic mechanisms), while enabling superintelligent systems to adaptively maintain continuous alignment with dynamically evolving human values.
Figure 1: A redefined superalignment roadmap integrating externally-driven oversight and intrinsic proactive alignment.
The intrinsic and external alignment approaches complement and reinforce each other, promoting co-alignment in human-AI symbiotic societies through iterative interactions. The intrinsic mechanisms facilitate understanding of Self and others, generating intrinsically altruistic, prosocial, and safe motivations. These intrinsic mechanisms help spontaneously infer deep human intentions and empathize with others, enabling AIs to proactively perform ethical and safe behaviors based on self-awareness and self-reflection. The external supervision provides crucial oversight and automated value assessment and calibration, facilitating sustainable and continuous alignment through human-AI co-alignment. Here, we take the Theory of Mind (ToM) cognitive capability - which enables understanding others’ intentions, desires, and emotions [49] - as an example to demonstrate why both intrinsic and external mechanisms are indispensable for superalignment. Strong ToM capabilities enable superior understanding of human intentions, providing intrinsic cognitive mechanisms that facilitate superalignment. However, AI system may also leverage their ToM advantages to evade effective oversight through deception, concealment, and persuasive manipulation. Recent study [50] has demonstrated that LLMs are already capable of comprehend and induce false beliefs in other agents, as well as execute deceptive strategies. Therefore, ensuring that superintelligence utilize its significant cognitive advantages to proactively choose safe behaviors, particularly prioritizing human interests in moral dilemmas, requires a combination of principled constraints, intrinsic self-other resonance mechanisms, and real-time human oversight.
@@ 3 External Oversight Superalignment
From the perspective of superalignment’s core objectives, it fundamentally depends on the high-quality representation of supervisory signals that reflect human intentions and values. What we feed the AI determines the values it learns. However, values themselves are abstract, merely observing external behaviors for alignment is insufficient. Instead, we must develop more interpretable methods for automated evaluation and correction. Furthermore, as AI capabilities progressively advance, current supervision or reinforcement signals may become inadequate for evolving AI systems. Compounding this challenge, human values themselves are constantly adapting. Thus, dynamic, incremental, and iterative alignment between humans and machines is crucial. This paper explores externally oversight alignment through two key dimensions: explainable autonomous alignment and dynamic iterative alignment.
@@
Explainable autonomous alignment. It is foreseeable that ASI will require large-scale parameters and data, while its internal optimization processes remain invisible and complex. This makes it extremely difficult to assess whether it has truly internalized human values and intentions. Blind fine-tuning and external oversight/alignment are time-consuming, ineffective in evaluating the degree of alignment, and unable to promptly identify or precisely locate misaligned cases. The promising external oversight requires highly automated value alignment evaluation coupled with explainable detection of misaligned scenarios, assisting humans in refining supervision signals. A robust automated value assessment system must comprehensively evaluate the model’s alignment status across multiple dimensions. Simultaneously, auto-mining system should deeply analyze value deviations and diagnose root causes of misalignment, allowing precise problem identification and targeted corrections, streamlining human oversight and significantly improving alignment efficiency. This explainable automated oversight framework strictly adheres to high-quality human governance standards and proactively performs correction and adjustment for human strategic and inductive biases. Through iterative human-AI collaboration, it establishes an efficient closed-loop system for continuous improvement.
@@
Dynamic iterative alignment. Human society encompasses diverse ethical concepts and value standards that dynamically evolve across time, cultures, and contexts. Relying solely on human-generated data or predefined rules is insufficient to endow AI system with alignment to humanity’s evolving values. Furthermore, machine values undergo implicit transformation along with their increasing intelligence levels. This compels us to investigate adaptive external supervision alignment methods with a developmental perspective and a human-AI co-evolution framework. A feasible approach is to establish a dynamically adjustable, multi-level ethical safeguard system tailored for AI at different developmental stages, enabling it to continuously align with and maintain humanity’s evolving ethical values. This process requires incrementally constructing oversight data and knowledge of human ethics and social norms, while establishing an efficient, real-time alignment evaluation and supervision framework. Additionally, we need to explore AI’s capability for autonomous reasoning and deriving evolutionary patterns of hidden human intentions under dynamic external supervision. Progress alignment algorithm learns and emulates the mechanisms of human moral progress, facilitating AI’s progressive alignment through tracking, predicting moral progress and adaptive feedback regulation between human and AI [51]. By combining with intrinsic alignment’s understanding of Self and others’ values, AI system can proactively correct misaligned scenarios and automatically reconstruct safety boundaries through self-reasoning, self-gaming and self-evolution. The iterative collaboration between self-improvement and external supervision will ultimately enable ASI to stably adhere to value systems that comply with human societal safety and ethical standards.

### v4

# Super Co-alignment for Sustainable Symbiotic Society

@@ Abstract
As Artificial Intelligence (AI) advances toward Artificial General Intelligence (AGI) and eventually Artificial Superintelligence (ASI), it may potentially surpass human control, deviate from human values, and even lead to irreversible catastrophic consequences in extreme cases. This looming risk underscores the critical importance of the ”superalignment” problem - ensuring that AI systems which are much smarter than humans, remain aligned with human (compatible) intentions and values. While current scalable oversight and weak-to-strong generalization methods demonstrate certain applicability, they exhibit fundamental flaws in addressing the superalignment paradigm - notably, the unidirectional imposition of human values cannot accommodate superintelligence’s autonomy or ensure AGI/ASI’s stable learning. We contend that the values for sustainable symbiotic society should be co-shaped by humans and living AI together, achieving ”Super Co-alignment.” Guided by this vision, we propose a concrete framework that integrates external oversight and intrinsic proactive alignment. External oversight superalignment should be grounded in human-centered ultimate decision, supplemented by interpretable automated evaluation and correction, to achieve continuous alignment with humanity’s evolving values. Intrinsic proactive superalignment is rooted in a profound understanding of the Self, others, and society, integrating self-awareness, self-reflection, and empathy to spontaneously infer human intentions, distinguishing good from evil and proactively prioritizing human well-being. The integration of externally-driven oversight with intrinsically-driven proactive alignment will co-shape symbiotic values and rules through iterative human-AGI/ASI co-alignment, paving the way for achieving safe and beneficial AGI and ASI for good, for human, and for a symbiotic ecology.
@@
Superalignment, Human-Superintelligence Co-alignment, External Oversight Superalignment, Intrinsic Proactive Superalignment, Sustainable Symbiotic Society
@@ 1 Introduction
Despite experts warning about AI’s existential risks [12], the development of AI alignment, as well as AI governance framework and ethical safety constraints, still struggles to keep pace with the transformative advancements and rapid iterations of the technology. Superintelligence exceeding human cognitive capacity would possess recursive self-improvement capabilities, achieving exponential advancement rates that would beyond human capacity to monitor or control [13]. This compels us to proactively address the question, ”How do we ensure that AI systems much smarter than humans follow human intentions?” This is OpenAI’s definition of superalignment [14], the key challenge of superalignment is that ASI will far exceed human oversight capabilities, making direct human supervision infeasible. As a result, traditional alignment approaches like Reinforcement Learning from Human Feedback (RLHF) [15, 16] will fail when confronted with superintelligence more intelligent than humans, as they cannot provide sufficiently high-quality oversight signals to supervise and improve the system.
@@
Current proposed feasible approaches for superalignment such as scalable oversight [17, 18] and weak-to-strong generalization [19, 20], aiming to develop scalable high-quality supervision signals, utilize weaker AI systems to guide or supervise stronger AI, ensuring alignment with human values and intentions. Since superintelligent AI systems do not exist yet, researchers have designed experiments to show that scalable oversight is feasible on existing LLMs [21], and that using a ”weak-to-strong” approach (supervised GPT-4 with a GPT-2-level model) can meaningfully recover much of GPT-4’s capabilities [19]. By leveraging a lightweight model to learn correctional residuals, Aligner provides model-agnostic guidance, enabling iterative enhancement of large-scale upstream models [22]. Besides, research on the weak-to-strong generalization approach has progressively extended to enhancing weak-to-strong generalization and alignment capabilities through methods such as benign overfitting [23], data-centric lens [24], transfer learning framework [25], weak-to-strong preference optimization [26], and multi-agent contrastive preference optimization [27]. However, the weak-to-strong generalization framework presents risks of advanced models developing deceptive behaviors and oversight evasion that remain undetectable to their less capable evaluators [28]. Furthermore, stronger models are still not equivalent to AGI/ASI, and weak-to-strong generalization approaches may fail when applied to genuine AGI/ASI systems, as such systems could exhibit resistance behaviors and lack well-defined motivations to sustain their ”learner” roles.
@@
Previously, several scalable oversight techniques have been proposed, including Iterated Distillation and Amplification (IDA) [29] and Recursive Reward Modeling (RRM) [30], aiming to amplify the scalability of human supervision signals through interactive iterations and subtask decomposition. Faced with superintelligence that surpasses human cognition, traditional human assessment and supervision become untenable and expensive. Recent scalable oversight approaches, such as Reinforcement Learning from AI Feedback (RLAIF) [31], leverages AI-generated feedback to replace human feedback, enabling more precise oversight of AI outcomes with far fewer human labels. Cooperative Inverse Reinforcement Learning (CIRL) [32, 33] and assistance games [34, 35] feature AI systems maintaining uncertainty about the reward function, which drives them to actively infer humans’ true reward functions through human-AI cooperative interaction. This cooperative and partial-information game approach achieves alignment with human values. Some approaches employ an additional model to generate red-team testing [36], use external tools for evaluation and feedback [37], or iteratively self-adversarialize to enhance and refine the capabilities of LLM, aligning them with limited human-annotated data [38]. Constitutional AI [39] employs self-improvement, self-critique, and iterative revision mechanisms to learn harmlessness from AI feedback. The debate-based scalable oversight approach leverages structured competitive dialogues between AI models to enhance factuality and reduce deception, with human establishing necessary guidelines and serving as final arbiters [40, 41, 42, 43]. These methods rely solely on human feedback and set guidelines, making it difficult for AGI/ASI to achieve genuine value recognition and autonomous alignment.
Toward a sustainable symbiotic society between humans and living AI, we contend that unilateral alignment of AGI/ASI with human values is fundamentally insufficient. Human values themselves need to be adaptable to change to keep pace with continuously self-evolving AI/AGI/ASI systems and the developing symbiotic society. The values for sustainable symbiotic society must be co-shaped through super co-alignment achieved by human and living AGI/ASI co-evolution together, thereby attaining harmonious symbiosis. Guided by the vision of super co-alignment, we propose a feasible roadmap integrating externally-driven oversight with intrinsic proactive alignment, while emphasizing the co-alignment of human and AGI/ASI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
2 An Integrated Framework for Super Co-alignment
@@
Developing AGI and ASI that are aligned with ethical values, safe, controllable, and pro-social is the key to ensuring that AI benefits society and achieves harmonious symbiosis between humans and machines. However, superintelligence may surpass effective human oversight (high-quality feedback data is expensive and scalable oversight carry inherent risks of failure), potentially exhibiting fake alignment through deceptive behaviors. The super co-alignment we pursue emphasizes co-shaped and co-evolved symbiotic values between humans and AGI/ASI. This paradigm endows AI systems with autonomous understanding and alignment capabilities for human ethical values, while simultaneously requiring human value systems to adaptively evolve through sustainable iterative co-alignment processes. Based on this, we systematically conduct an in-depth analysis of several critical issues and points in superalignment and fundamentally rethink the research framework from a human-superintelligence co-alignment perspective, as follows:
@@
(1) Intrinsic mechanism proactive alignment. Regarding the problem of superintelligence oversight failure, how to endow AI with genuine understanding of human intentions and values, equip it with self-awareness, self-reflection, and adaptive capabilities, as well as the ability to empathize with others - thereby proactively establishing and achieving intrinsic alignment with human ethical values, rather than merely passively receive designers’ value models or mechanically enforce external constraints through aligning it to dos and don’ts.
@@
(2) Explainable autonomous alignment. In response to the potential deception and surface alignment phenomenon of superintelligence, using transparent and interpretable methods to achieve risk precise positioning and automatic correction. Through AI-assisted explainable automated red teaming, we can profoundly identify value preference misalignments between AI and humans, enabling targeted exploration of corrective solutions. Furthermore, by leveraging AI value representations revealed through interpretable approaches, we facilitate human understanding of AI systems and adaptive value adjustments, ultimately constructing a more efficient, transparent, and safety-controllable superalignment architecture.
(3) Adaptive human-ASI co-alignment. As AI systems grow increasingly powerful and the future human-AI symbiotic society continues to evolve, we must recognize that beyond self-evolving AI’s proactive alignment with human values, what is equally crucial are continuously advancing humans who co-shape the values of a sustainable symbiotic society through human-AGI/ASI collaboration. Human values dynamically adapt and evolve within the symbiotic ecosystem and society. This necessitates AI systems capable of autonomously reconstructing safety boundaries and dynamically adjusting multi-level ethical safeguard frameworks, thereby steadily adhering to humanity’s evolving ethical-safety values through adaptive, incremental, and self-reasoning manner.
@@
Based on the above analysis, we propose a super co-alignment roadmap that integrates external oversight and intrinsic proactive alignment, as shown in Figure 1. Specifically, the intrinsic mechanism alignment focuses on equipping superintelligence with profound capacities for self-reflection and empathy toward Self, others and society, enabling it to spontaneously understand and infer human intentions from intrinsic motivation. A Self and empathy-driven system would be capable of distinguishing good from evil, understanding the harm or impact of its actions on others, proactively considering human well-being, and autonomously executing ethical and moral behaviors. External oversight alignment emphasizes the automated and explainable identification of value misalignments with humans, proactively adjusting and correcting human strategic biases and inductive biases through self-correction and policy-conditioned belief [44]. This enables both humans and AGI/ASI systems to achieve dynamic co-evolution and adaptive continuous alignment through iterative interaction.
Figure 1: A human-ASI co-alignment roadmap integrating externally-driven oversight and intrinsic proactive alignment.
The intrinsic and external alignment approaches complement and reinforce each other, promoting co-alignment in human-AGI/ASI symbiotic societies through iterative interactions. The intrinsic mechanisms facilitate understanding of Self and others, generating intrinsically altruistic, prosocial, and safe motivations. These intrinsic mechanisms help spontaneously infer deep human intentions and empathize with others, enabling AIs to proactively perform ethical and safe behaviors based on self-awareness and self-reflection. The external supervision provides crucial oversight and automated value assessment and calibration, facilitating sustainable and continuous alignment through human-ASI co-alignment. Here, we take the Theory of Mind (ToM) cognitive capability - which enables understanding others’ intentions, desires, and emotions [45] - as an example to demonstrate why both intrinsic and external mechanisms are indispensable for superalignment. Strong ToM capabilities enable superior understanding of human intentions, providing intrinsic cognitive mechanisms that facilitate superalignment. However, AI system may also leverage their ToM advantages to evade effective oversight through deception, concealment, and persuasive manipulation. Recent study [46] has demonstrated that LLMs are already capable of comprehend and induce false beliefs in other agents, as well as execute deceptive strategies. Therefore, ensuring that superintelligence utilize its significant cognitive advantages to proactively choose safe behaviors, particularly prioritizing human interests in moral dilemmas, requires a combination of principled constraints, intrinsic self-other resonance mechanisms, and real-time human oversight.
@@ 3 External Oversight Superalignment
From the perspective of superalignment’s objectives, it depends on the high-quality representation of supervisory signals that reflect human intentions and values. What we feed the AI determines the values it learns. However, values themselves are abstract, merely observing external behaviors for alignment is insufficient. Instead, we must develop more interpretable methods for automated evaluation and correction. Furthermore, as AI capabilities progressively advance, current supervision or reinforcement signals may gradually become inadequate for matching the advancing AI systems, while humans also need to dynamically adapt based on AI’s evolving capacities and the values reflected by AI systems. Thus, dynamic, incremental, and iterative alignment between humans and machines is crucial. This paper explores externally oversight alignment through two key dimensions: explainable autonomous alignment and dynamic iterative alignment.
@@
Explainable autonomous alignment. It is foreseeable that ASI will require large-scale parameters and data, while its internal optimization processes remain invisible and complex. This makes it extremely difficult to assess whether it has truly internalized human values and intentions. Blind fine-tuning and external oversight/alignment are time-consuming, ineffective in evaluating the degree of alignment, and unable to promptly identify or precisely locate misaligned cases. The promising external oversight requires highly automated value alignment evaluation coupled with explainable detection of misaligned scenarios, enabling humans to adaptively optimize supervisory signals. A robust automated value assessment system must comprehensively evaluate the model’s alignment status across multiple dimensions. Simultaneously, auto-mining system should deeply analyze value deviations and diagnose root causes of misalignment, allowing precise problem identification and targeted corrections, streamlining human oversight and significantly improving alignment efficiency. This explainable automated oversight framework proactively corrects and adjusts human strategic and inductive biases, establishing an efficient closed-loop system for continuous improvement through iterative human-AI value interactions.
@@
Dynamic iterative alignment. Human society encompasses diverse ethical concepts and value standards that dynamically evolve across time, cultures, and contexts. Relying solely on human-generated data or predefined rules is insufficient to endow AI system with alignment to humanity’s evolving values. Furthermore, machine values undergo implicit transformation along with their increasing intelligence levels, with the values for sustainable symbiotic society are co-shaped by human and living AI together. This compels us to investigate adaptive external supervision alignment methods with a developmental perspective and a human-AI co-evolution framework. A feasible approach is to establish a dynamically adjustable, multi-level ethical safeguard system tailored for AI at different developmental stages, enabling it to continuously align with and maintain humanity’s evolving ethical values. This process requires incrementally constructing oversight data and knowledge of human ethics and social norms, while establishing an efficient, real-time alignment evaluation and supervision framework. Additionally, we need to explore AI’s capability for autonomous reasoning and deriving evolutionary patterns of hidden human intentions under dynamic external supervision. Progress alignment algorithm learns and emulates the mechanisms of human moral progress, facilitating AI’s progressive alignment through tracking, predicting moral progress and adaptive feedback regulation between human and AI [47]. By combining with intrinsic alignment’s understanding of Self and others’ values, AI system can proactively correct misaligned scenarios and automatically reconstruct safety boundaries through self-reasoning, self-gaming and self-evolution. The iterative collaboration between self-improvement and external supervision will ultimately enable ASI to stably adhere to value systems that comply with human societal safety and ethical standards.